Find the peak to peak voltage, peak voltage, RMS voltage if the sinusoidal wave in CRO has peak to peak difference of 4.4 divisions and the scope scale is set to 0.1 mv/div
1
Expert's answer
2021-10-24T01:17:45-0400

Find the peak to peak voltage, peak voltage, RMS voltage if the sinusoidal wave in CRO has peak to peak difference of 4.4 divisions and the scope scale is set to 0.1 mv/div


Given, Scope Scale = 0.1 mv/div.

Peak to peak difference = 4.4 divisions

Therefore, the peak division 4.4/2 = 2.2 division


Hence peak voltage = 2.2 * 0.1 = 0.22 mv


RMS Voltage = "RMS Voltage = V_{rms} = \frac{V_{peak}}{sqrt(2)} = \frac{0.22}{1.414} = 0.155 mV"